OpenStudy (anonymous):

A photograph with a width of 3 inches and a length of 5 inches will be enlarged so that the length is 40 inches. What will the width equal? A) 20 inches B) 24 inches C) 88 inches D) 120 inches

OpenStudy (anonymous):

24 inches

OpenStudy (anonymous):

because 3/5 = x/40 then (3/5)40 = x 3*8=X 24=x